OpenStudy (anonymous):

solve the differential equation y'=x+5y

OpenStudy (lgbasallote):

this is linear right?

OpenStudy (anonymous):

for whatever reason I wanna take the derivative but that's not what I'm supposed to do . Yes linear

OpenStudy (lgbasallote):

\[\LARGE \frac{dy}{dx} - 5y = x\] use integrating factor \[e^{\int -5dx} = e^{-5x}\] then do that \[y(I.F.) = \int Q(I.F.)dx\] thingy \[\large y(e^{-5x}) = \int xe^{-5x}dx\] where did you get confused?

OpenStudy (anonymous):

ok I have never done this before

OpenStudy (anonymous):

e?

OpenStudy (anonymous):

ok lets see what's going on... we have \[e^{\int -5dx}\] and \[e^{ -5x}\]

OpenStudy (lgbasallote):

oh you havent? it's just \[\frac{dy}{dx} + Py = Q\] where P and Q are functions of x then you solve for integrating factor \[e^{\int Pdx}\] once you have solved the integrating factor, you use the equation \[y(I.F.) = \int Q(I.F.)dx\] where I.F. means integrating factor

OpenStudy (lgbasallote):

are you not familiar with integrating factors in linear DE?

OpenStudy (anonymous):

I'm not sure, P in this case is 5x

OpenStudy (lgbasallote):

no it's just 5

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

wait it's -5

OpenStudy (anonymous):

how about Q 1?

OpenStudy (lgbasallote):

Q would be x \[\large \frac{dy}{dx} + Py = Q \implies \frac{dy}{dx} - 5y = x\]

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):